3) a) If the corporate tax rate is 30% and the individual tax rate is 10% (and these are the only taxes), a corporation with $20,000 of operating income, what is the after-tax disposable income the corporation generated for its owners? The tax rate is 5% between $0-$10,000 and 12% between $10,001 _$20,000. If you earn $12,000, then what is the average tax rate and the marginal tax rate? b) 4) a) You decide that your family would be comfortable living on an annual income of $150,000, growing at 4% per year. Youd also like to continue generating this cash flow for your descendants, forever. With interest rates of 8%, how much wealth would you need today to provide this income starting with $150,000 one year from now? b) Suppose all income is taxed at 20%. How would this (if at all) change your answer to a)? Assume that you still feel 4% growth and $150,000 (after-tax) income is sufficient. Show a calculation and briefly explain
Answer #1

If operating income is $20000 and corporate tax is 30% then Profit after tax is 20000(1-0.3)=14000

Now disposable income will be $14000(1-0.1)=$12600

Income generated by corporation to owners would be $12600

Ans B)

Average tax rate would be calculated as below

Tax in first bracket =10000*0.05=$500

Tax in second bracket =(12000-10000)*0.12=240

Average tax rate is 500+240/12000=740/12000=6.167%